Lots going on with AAVSOnet! The major software efforts have been: to interface with two roof controllers, making automated operation of BSM_NM and BSM_TX more reliable; working on the dual filter wheel plugin for MaximDL; and continuing the port of the HQ pipeline to Python. George Silvis and Cliff Kotnick have been putting in some long volunteer hours on the software, with very positive results! Ken Menzies has been instrumental in creating new plans for the many recent proposals, including some very complex ones in support of HST targets. Dave has been interfacing with HQ and the Board, and keeping everyone on track.
BSM_Berry. This system is currently using the AT72 refractors, along with one of the ZWO ASI183mm-pro CMOS cameras, and is fully operational. The E180 and Paramount will be installed in the future.
BSM_Hamren. This system has been shuttered at Cape Cod. Gary Walker has packaged it for shipment to Hawaii. Bill Wiecking at the Hawai'i Preparatory Academy (HPA) will be the new host of BSM_Hamren. We hope to have it on-line within the next month. Our hats off to Mike Linnolt for his multi-year hosting of this system!
BSM_NH2. Fully operational. We're still experimenting with the dual filter wheel configuration, so automated switching between the ~32 positions is not yet available.
BSM_NM. Fully operational. George Silvis has written scripts to interface with the Backyard Observatory roof controller, so that images are not taken with a closed roof.
BSM_S. Peter Nelson has installed the ZWO ASI183mm-pro CMOS camera on the AT72 refractor. He is still working out some bugs regarding a stable transfer connection, but it should be on-line soon.
BSM_TX. Mostly operational. George Silvis has written scripts to interface with the roof controller, so that images are not taken with a closed roof. There appears to be some slippage in the focuser, so that some images are sharply focused and some are not. We're investigating this issue.
OC61. Fully operational. There are some instances of the mount controller pointing the telescope horizontal, a situation that has to be manually solved. The new QHY600 camera and new mount controller will be shipped to Mt. John soon.
SRO. The camera may have failed. We're investigating the problem. This STL6303 camera has been in continuous operation for many years, so problems are not unexpected. It should be replaced.
TMO61. This system is again operational, using the old QSI-583 camera. A QHY600 camera has been purchased, and when all of the filters arrive, will be installed.
Sad to hear SRO is out of action. Will observations planned for SRO under Proposal 153 be moved to another instrument? I was expecting to requst resumption of observations of IX Dra on Jan 1st to capture its January superoutburst.
My last image from BSM NM(2) was dated 10 Dec for Proposal 163 monitoring RZ Lmi and Er Uma. Has the instrument been clouded out or has there been a tech issue?